Given a geometric sequence whose third term 3/4 and seventh term is 243/64. Find the fifth term of the sequence

3 243 18. and seventh term is Given a geometric sequence whose third term 4 64 Find the fifth term of the sequence.
Answer #1

Aug know general form ( nth term) of geometric sequence cohere i.e. is first term Th=ar(n-1) G r = Common radio a Tn {na Thol3 X 3 X 3x3 ll 1 X2 X 2 X 2 + 34 29 que 04 34*14 24xYg 2. . mor Now ar2 = plus (equation ₂ 2 a13 2 :) ах 9 4 - 3 3 4 g3 a= wh